5:12pm: We have been made aware of an issue whereby a small number of customers are experiencing issues accessing their websites. The issue was first raised at 4:57pm on Friday 28th July 2017. Our engineers are looking into the issue and further updates will be posted at 5:57pm.
5:49pm: ESP engineers are onsite at our data centre working on the issue. A number of customers have had their service fully restored but a small number are yet to be resolved. A further update will be due at 6:49pm
6:34pm: ESP engineers are still working on the issue. It has been identified that their were some hardware issues with some UPS devices at the data centre. Our engineers are working hard to find a work around until replacement hardware can be sourced on Monday. A further update will not be due until 9pm.
9:07pm: Engineers are still working on the issue. We have no update at this time. A further update will be due at 11:07pm.
11:33pm: Apologies for the late update. All our engineers have been working hard to resolve the issues and we are wrapping up for the day. Work will be starting from 9:30am tomorrow. A further update will be provided before 11:30am (20/07/2017)
11:25am: Engineers have resolved the issues and service is now fully restored. We are going to be investigating the root cause of these issues over then next few days.